OpenStudy (anonymous):

whats is square root of 96 in simplest radical form

OpenStudy (amistre64):

what have you tried?

OpenStudy (amistre64):

a good strategy is to see if 96 is divisible by a perfect square like; 4,9,25,36,etc

OpenStudy (amistre64):

or start dividing it by familiar numbers and see what pops up

OpenStudy (anonymous):

96=12 x 8 square root of 12 = \[\sqrt{2\times2\times3}\] So it will be \[2\sqrt{3}\] And for 8, \[\sqrt{2\times2\times2}\] So It is \[2\sqrt{2}\] Multiplying both, \[2\times2 \sqrt{2\times3}\] \[4\sqrt{6}\] is the answer.

OpenStudy (amistre64):

i put sqrt(8) = 4sqrt(2) on my test the other day :/

OpenStudy (anonymous):

Hahaha we all make mistakes sometimes man.